The Gauteng traffic department records show that 25% of all drivers wear seatbelts. In a random sample of 400 cars stopped at a roadblock in Gauteng, 152 of the drivers were wearing seatbelts. What is the probability that at most 18% of the drivers in Gauteng wear seatbelts?



1
Expert's answer
2021-10-04T16:14:40-0400

"p=0.25 \\
n=400 \\
\hat{p} \sim (p, \frac{p(1-p)}{n}) \\
Z = \frac{\hat{p} -p}{\sqrt{\frac{p(1-p)}{n}}} \\
P(\hat{p} ≤0.18) = P(\frac{\hat{p} -p}{\sqrt{\frac{p(1-p)}{n}}} ≤ \frac{0.18-0.25}{\sqrt{\frac{0.25 \times 0.75}{400}}}) \\
= P(Z≤ -3.233) \\
= 0.0006"